¡GoConqr estrena diseño! Cuéntanos qué te parece.

Apuntes

Stastics

kia_stevens Inglés 1 página

Contenido del apunte (1)

Página 1

1.Probability;

Probability is a measure of how likely an event is

You can use numbers instead of words

An outcome is what might happen

Think about all possible outcomes

You can find probabilities of events using a formula

P(event)=Number of outcomes in the (successful/unsuccessful)/ total number of outcomes

The probabilities of all outcomes add up to 1

Probabilities are used to asses risk

2.Sample Space Diagrams;

A sample space is a list of all possible outcomes

Use a sample space diagram to list all the outcomes

A venn diagram shows a sample space too

You can use venn diagrams to work out probabilities

Two way tables can be used to find probabilities too= two pieces of data

3.Relative Frequency

Its an Estimated probability

Relative frequency is the frequency relative to the total number of trials

Probability of something happening= Number of times it happened/ Number of trials

The more trials there are, the more accurate the probability will be

4.Expected and Actual Frequencies;

Expected Frequency is what's likely to happen

Once you know the probability you can predict how many times it is expected to happen

Expected Frequency=Number of trials X the probability of the outcome happening

Results might be different from expected frequencies, compare results with actual results

5.Probability Laws;

Mutually Exclusive=Events cant happen together, cant happen at the same time

P(A or B)= P(A) + P(B)

The general addition law is for Non-Mutually Exclusive Events, when 2 events happen together

P(A or B)= P(A) + P(B) -P(A and B)

Exhaustive events are at least one of them must happen.

If you add together the probabilities of exhaustive mutually exclusive events you always get 1

Independent events are unconnected, they have no effect on each other

The rule of independent events saves a lot of counting, multiply the probabilities of each event together

P(A and B)= P(A) x P(B)

You can find out the probability of at least 1

P(at least 1) = 1- P(none)

6.Tree Diagrams;

You can use tree diagrams for probability questions

Always multiply along the branches

Probabilities must always add up to 1

Add up relevant results

7.Conditional Probability;

When one event affects another

You dont always need the whole sample space, sometimes only concentrate on one bit

General law for combining events

P(A and B)= P(A) x P(B given that A occurs)

Use tree diagrams to answer questions involving conditional probabilities

Conditional Probabilities= P(B given that A occurs)= P(A and B) / P(A)

8.Discrete Probability Distributions;

Discrete Uniform distribution are equal chances, You get Discrete Uniform Distributions when the probability of each outcome is EQUAL

Binomial Distribution are 2 mutually exclusive outcomes. E.g, Heads or tails

New Page

Información

Descripción
Apunte sobre Stastics, creado por kia_stevens el 29/12/2014.
Creado por
kia_stevens
Visitas
14
Valoración

Escríbenos

¿Sobre qué es tu mensaje?

¿Qué te parece el nuevo diseño?

¿Te gusta?